OpenStudy (anonymous):

7 sin2x - 14 sin x + 2 = -5 find all solutions

OpenStudy (anonymous):

cosx=-1

OpenStudy (anonymous):

thats not even english to me.

OpenStudy (anonymous):

ok

OpenStudy (hba):

@hartnn

OpenStudy (anonymous):

7sin²(x) - 14sin(x) + 2 = -5 7sin²(x) - 14sin(x) + 7 = 0 7(sin²(x) - 2sin(x) + 1) = 0 sin²(x) - 2sin(x) + 1 = 0 (sin(x) - 1)² = 0 sin(x) - 1 = 0 sin(x) = 1 x = 90°

OpenStudy (anonymous):

would it be that?

hartnn (hartnn):

is it sin2x or \(sin^2 x\) ??

OpenStudy (anonymous):

sin squared x

hartnn (hartnn):

then sin x=1 is correct. but do u need solutions in the range of 0 to 180 only or all solutions ?

OpenStudy (anonymous):

no idea. the question say find all solutions to the equation

hartnn (hartnn):

because sin (360+90) also =1

hartnn (hartnn):

yes, all solutions means u don't have only 1 solution, u have infinite

OpenStudy (anonymous):

alright

hartnn (hartnn):

sin(x)=1 sin x = sin (90+360n) (n=1,2,3,....) so x= 90+360n are all the values of x with n=1,2,3,....`

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

yeah. is that the answer

hartnn (hartnn):

yes.

OpenStudy (anonymous):

you sure. i can wrrite that & ill be correct?

OpenStudy (anonymous):

if i wrote exactly what you just put down ill be correct?

hartnn (hartnn):

idk how your teacher / system demands the answer, but any of these 2 answers are complete and correct: 1) x= 90+360n , n=1,2,3,..... 2) x=90,450,810,..

OpenStudy (hba):

Whats The Range ? I think It Would Be 0 to 180

OpenStudy (anonymous):

thanks